Logbooks Lab Maintenance Evaporator_1 Evaporator_2 Laser cutter Target Production Test-Stand RH-ISAC RH-Cyclotron RH-Meson Hall RH-Beamlines RH-ARIEL
  RH-ISAC, Page 132 of 140  Not logged in ELOG logo
ID Date Author Category Type Module Target/Number Subject Text Attachments
  182   Tuesday, September 18, 2012, 13:47 chad fisherSouth Hot-CellStandard OperationTM1Nb#7Nb #7 install Cover BaselineNb #7 was installed and tested today.
Inside face of the containment box
cover was cleaned with methanol and a baseline
 IMG_3191.JPGIMG_3192.JPG 
  181   Tuesday, September 18, 2012, 13:41 David WangSouth Hot-CellStandard OperationTM1Nb#7leak check heat shield circuitChad  and i did a leak check on TM1 heat
shield circuit with new Nb#7 target  installed.
It took about 5 mins pump down  the circuit
  
  180   Tuesday, September 18, 2012, 09:13 Travis CaveSouth Hot-CellStandard OperationTM1no targetmodule moveTM#1 moved from the south west silo to
the south hot cell, move went well.
  
  179   Friday, September 14, 2012, 10:45 Travis CaveSpent Target VaultStandard Operation UCx#4Spent target moveUCx#4 has been moved from the south hot
cell to the spent target vault. The target
was placed in pail #99 and now resides in
 Vault_Storage_September_14_2012.pdf 
  178   Thursday, September 13, 2012, 13:27 Travis CaveSouth Hot-CellStandard OperationTM1no targetmodule moveTM#1 has been moved from the south hot
cell to the south west silo. Move went well.
  
  177   Thursday, September 13, 2012, 11:03 Travis CaveSouth Hot-CellStandard OperationTM1UCx#4module moveTM#1 with UCx#4 has been moved from the
west target station to the south hot cell.
Move went well.
  
  176   Wednesday, September 12, 2012, 11:59 David WangITWStandard OperationTM1UCxTM1 disconnected in ITWI removed 5 shielding blocks from ITW to
block storage area, and disconnected TM1
in ITW. After UCx run,  the radioactive
  
  175   Wednesday, September 12, 2012, 07:00 David WangITEStandard OperationTM4NiO #1ITM4 in ITE connected and pumping downTM4 had been connected in ITE. Six turbo
pump had been started . thepumpming diwn
is very good so far. The water interlock
  
  174   Friday, September 07, 2012, 10:37 Travis CaveConditioning StationStandard OperationTM4NiO#1Electrical checkElectrical check of the TM#4 with NiO#1
was done again today. see attached file for
results.
 TM#4_NiO#1a.PDF 
  173   Friday, September 07, 2012, 09:47 Travis CaveSpent Target VaultStandard Operation UO#3Spent target moveThe spent UO#3 was moved from the south
hot cell to the storage vault. It was in
pail 98 and placed in spot 1A of the vault,
 Vault_Storage_September_6_2012.pdf 
  172   Friday, September 07, 2012, 08:54 David WangConditioning StationStandard OperationTM4NiO #1TM4 NiO #1 target heat shield helium pressure check is done.leak detector pumped down to 0.0xE-9 atm.cc/sec,
and 0.0xE-4 torr. All high vacuum gauge readings
at CS and TM4 are at low -6 range from 2.2xE-6torr
  
  171   Wednesday, September 05, 2012, 14:40 Travis CaveConditioning StationStandard OperationTM4NiO#1Electrical checkSee attached PDF for details. 60kV Common
to the target is low at 343k but I've been
told will work. Test will be done again on
 TM#4_NiO#1.PDF 
  170   Wednesday, September 05, 2012, 14:36 chad fisherSouth Hot-CellStandard OperationTM4NiO #1UOx removal NiO installationUOx was removed from TM4 yesterday and
NiO#1 installed.
Video was taken.
  
  169   Wednesday, September 05, 2012, 11:55 Travis CaveConditioning StationStandard OperationTM4NiO#1module moveTM#4 with NiO#1 has been moved from the
south hot cell to the conditioning station.
Move went well.
  
  168   Tuesday, September 04, 2012, 14:34 David WangSouth Hot-CellStandard OperationTM4NiO #1TM4 heat shield circuit leak check with NiO#1 targetChad and i did a leak check on TM4 heat
shield circuit with new NiO#1 target installed.
It took about 40 mins(much longer than usual)
  
  167   Tuesday, September 04, 2012, 11:11 Travis CaveSouth Hot-CellStandard OperationTM4UO#3module moveTM#4 with UO#3 has been moved from the
east target station to the south hot cell
for a target exchange. The move went well.
  
  166   Friday, August 31, 2012, 12:10 chad fisherSouth Hot-CellMaintenance  Hot Cell MaitenanceEntered south hot cell Wednesday and replaced
the pre-filter air filter as well as the
front two light bulbs. Also removed two bags
  
  165   Thursday, August 30, 2012, 13:57 David WangITEStandard OperationTM4 Disconnected TM4 in ITE.I disconnected TM 4 in ITE on AUG 29th
morning.The vacuum lines stay untouched and
they will be disconnected on SEP 04th.
  
  164   Thursday, August 23, 2012, 07:34 David WangITWRepair  Replaced broken air lines on ITW HARP5B cylinder.Prob guys replaced broken air lines on
ITW HARP 5B Cylinder.Several shielding blocks
had been moved out to make the cylinder accessible.
  
  163   Monday, August 20, 2012, 12:28 David WangITWStandard OperationTM1 I moved 5 shielding blocks from ITE to ITW,I moved 5 shielding blocks from ITE 
to ITW.ITW is configured to take proton beam
now.
  
ELOG V2.9.2-2455